# HyDeFuse: Provably Convergent Denoiser-Driven Hyperspectral Fusion

### v1

@@ 4.4 Discussion
After presenting our technical results, we now examine the connection between our analysis and related works [52, 42, 46, 49], which employ kernel denoisers to develop provably convergent PnP models. In [52, 42, 46], the kernel denoiser is designed as a proximal operator of a convex regularizer, allowing PnP to be linked to a convex optimization model. Standard results on proximal algorithms [72] are then used to establish objective convergence. However, in our case, the denoiser $\mathcal{D}$ is not necessarily self-adjoint, as $\mathcal{V}$ and $\mathcal{W}$ do not commute. This lack of symmetry implies that $\mathcal{D}$ cannot be expressed as the proximal operator of a convex regularizer, following a classical result by Moreau [73]. As a result, the approach in [52, 42, 46] is not directly applicable for HyDeFuse. Moreover, we establish a stronger result –– iterate convergence over objective convergence –– as well as the convergence rate. Our analysis instead builds on [49]. However, the loss function, the underlying vector space and operators, and the denoiser in HyDeFuse are different from the generic setup in [49]. As a result, the proofs from [49] do not directly apply, requiring us to refine and adapt the arguments.
@@ 5 Numerical Results
In this section, we validate our theoretical findings and demonstrate the empirical performance of HyDeFuse. Specifically, we provide numerical results on global convergence and the convergence rate, assess the proposed denoiser CasKD in comparison to existing denoisers, and benchmark HyDeFuse against various classical and state-of-the-art fusion methods.
@@ 5.1 Experimental Setup
We provide a detailed description of the dataset, the parameters of the forward model, the quality metrics used for evaluation, and the denoising and fusion methods used for comparison.
@@
Blur models. The Starck-Murtagh filter was applied as the blurring operator for the HS image [8, 52]. The downsampling factors used to generate $\mathbf{Y}_{h}$ were set to 4 for Pavia, 6 for Chikusei, and 3 for Paris. The spectral response $\mathbf{R}$ was computed following the approach in [8]. If the SNR for HS and MS images is mentioned, it is denoted as SNRh and SNRm, respectively. If not specified, it is assumed to be 20 dB for both HS and MS across all datasets. We also compare HyDeFuse with recent methods using a Gaussian point spread function with radius $7$ and standard deviation $2$ [74]. This is to demonstrate that HyDeFuse performs effectively under varying blurring models.
Compared methods. We compare our kernel denoiser CasKD with both traditional and recent denoising algorithms. Specifically, we compare CasKD against BM4D [75], DnCNN [39], FHDD-NLM [52], KBR [51], LLRT [50], FastHyMix [54], and NGMeet [53]. We compare our fusion method HyDeFuse with a range of classical and state-of-the-art methods: GSA [7], CNMF [17], GLP [13], MAPSMM [9], HySure [8], GTTN [74], and CTDF [18]. Furthermore, within the PnP-PGD framework, we compare with denoisers such as DnCNN [39], BM4D [75], and GMM [45].
Quality metrics. We employed standard quality metrics [8, 43, 74], including Peak Signal-to-Noise Ratio (PSNR), Root Mean Squared Error (RMSE), Spectral Angle Mapper (SAM), Erreur Relative Globale Adimensionnelle de Synthèse (ERGAS), and Universal Image Quality Index (UIQI). For reconstruction tasks, lower RMSE, SAM, and ERGAS values indicate better performance, while higher PSNR and UIQI values (maximum is $1$) suggest superior reconstruction quality.
@@ 5.2 Denoising Results
Methods PSNR (0.0196) UIQI (0.0196) PSNR (0.0784) UIQI (0.0784)
@@
Table 2: Denoising comparison on the Pavia dataset for Gaussian noise (standard deviation in brackets). The top two values are highlighted in bold, and the best value is underlined; this convention is followed in all the tables.
@@
Figure 4: Comparison of the proposed denoiser CasKD with existing denoisers on the Pavia dataset. The noise is Gaussian with standard deviation $0.019$. For visualization, only three bands are displayed.
Methods PSNR (0.0196) UIQI (0.0196) PSNR (0.0784) UIQI (0.0784)
Noisy 34.16 0.969 22.10 0.68
FHDD-NLM 44.77 0.977 35.52 0.969
BM4D 40.44 0.992 31.70 0.944
DnCNN 42.49 0.996 36.78 0.982
KBR 39.88 0.991 34.25 0.967
LLRT 41.08 0.997 35.82 0.981
FastHyMix $\mathbf{47.04}$ $\mathbf{0.998}$ 37.44 0.984
NGMeet $\mathbf{47.52}$ $\mathbf{0.998}$ $\mathbf{38.22}$ $\mathbf{0.986}$
CasKD 46.94 $\mathbf{0.998}$ $\mathbf{38.24}$ $\mathbf{0.987}$
Table 3: Same as Table 2 but for the Paris dataset.
The proposed denoiser CasKD is a combination of two kernel denoisers. Combining two denoisers does not always yield good results. However, we recall that the component denoisers act differently in CasKD –– they work on the inter- and intra-band correlations between patches. This is a possible explanation for the improved performance in Table 1. To further evaluate the performance of CasKD, we compare it with several existing denoisers: BM4D [75], DnCNN [39], FHDD-NLM [52], KBR [51], LLRT [50], FastHyMix [54], and NGMeet [53]. A visual comparison is shown in Figure 4. In Tables 2 and 3, we give a detailed comparison based on PSNR and UIQI values. We observe that CasKD performs competitively with these denoisers, and in some cases, it outperforms them.
@@ 5.3 Convergence Results
Figure 5: Convergence rate $\mu$ for different $\lambda$ appearing in the loss function (6). The parameters $\sigma_{1}$ and $\sigma_{2}$ are the standard deviations of the Gaussian (RBF) kernels used in $\mathcal{W}$ and $\mathcal{V}$.
We used the power method [67] to calculate the convergence rate of HyDeFuse. We see from (37) that the rate is given by $\mu$, the largest singular value of the linear operator $\mathcal{P}$. However, computing $\mu$ via SVD is impractical because $\mathcal{P}$ is too large to store. Instead, the power method estimates $\mu$ by iteratively applying $\mathcal{P}$ and its adjoint $\mathcal{P}^{*}$ [67], where it follows from (21) that $\mathcal{P}^{*}=\mathcal{G}\circ\mathcal{W}\circ\mathcal{V}$, since $\mathcal{V},\mathcal{W}$ and $\mathcal{G}$ are self-adjoint.
In Table 4, we use the Chikusei dataset to validate the contraction factor of $\mathcal{P}$ in HyDeFuse. This depends on the step size $\gamma$, the forward model parameters $\mathbf{A}$ and $\mathbf{R}$, and the denoiser $\mathcal{D}$. Specifically, we used the power method to compute $\mu$ for step sizes $0<\gamma<2/\beta$ and for different denoiser settings. We see that as predicted by Theorem 1, the contraction factor is <1<1. On the same dataset, we computed the contraction factor for different values of $\lambda$, which is used to balance the loss function. The results are shown in Figure 5. Again, we see that $\mu$ always remains below $1$. Additionally, we see that convergence is faster when $\lambda$ is neither too large nor too small, which is noteworthy as this range is also expected to yield optimal fusion performance.
Figure 6: The plot shows the successive differences between HyDeFuse iterates and the corresponding PSNR values for different step sizes $\gamma$ (expressed in units of $1/\beta$). We computed $\beta$ using the power method and found it to be $0.5998$ for this setup. We observe that convergence is achieved for $\gamma\in\left(0,2/\beta\right)$. However, as demonstrated with the example $\gamma=2.2/\beta$, convergence cannot be guaranteed when $\gamma$ falls outside this range. Note that the PSNR stabilizes within 10 iterations for optimally chosen $\gamma$.
Figure 7: Same as Figure 6 but for the Pavia dataset and using a Gaussian point spread function with radius $7$ and a standard deviation of 22 for the blur operator $\mathbf{B}$. In this case, $\beta=0.389$. Once again, empirical convergence is observed for $\gamma\in\left(0,2/\beta\right)$, while values of $\gamma$ outside this range lead to divergence. Notice that the PSNR stabilizes within 10 iterations for an optimal $\gamma$.
(a) ground truth
(b) $\mathbf{X}_{0}$ : all-ones
(c) $\mathbf{X}_{0}$ : all-zeros
(d) $\mathbf{X}_{0}$ : white noise
Figure 8: Fusion results on the Chikusei dataset. The reconstructions shown are obtained from HyDeFuse using different initializations $\mathbf{X}_{0}$. In all cases, the performance metrics remain identical: PSNR = 42.27, RMSE = 0.0076, ERGAS = 1.39, SAM = 1.79, and UIQI = 0.967.
@@
Table 4: Contraction factors for different step sizes and denoiser settings. The step size $\gamma$ is given in multiples of $1/\beta$; $\beta$ is computed using the power method and found to be $0.9757$. The denoiser settings $(\sigma_{1},\sigma_{2})$ represent the standard deviations of the Gaussian (RBF) kernels in $\mathcal{W}$ and $\mathcal{V}$, where $a=0.0196$, $b=0.039$, and $c=0.117$ on a scale of $[0,1]$.
@@
As further evidence of the contractive nature of the reconstruction operator $\mathcal{T}$, we conducted fusion experiments on the Chikusei dataset. The results show that, irrespective of the initialization, the iterations consistently converge to the same fixed point, confirming the uniqueness of the fixed points of the contractive operator $\mathcal{T}$.
Finally, we analyze the convergence of the iterates $\{\mathbf{X}_{k}\}$ generated by (9). A sample result on the Paris dataset for the Starck-Murtagh filter as the blurring operator is presented in Figure 6. A similar result is reported in Figure 7, but for Pavia and using a Gaussian point spread function as the blur model. As predicted by Theorem 1, the successive difference $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|$ approaches zero, which is confirmed by these figures. Importantly, observe that the PSNR stabilizes within 10–15 iterations. This pattern was consistently observed across all three datasets. An interesting finding is that when the step size $\gamma$ exceeds the admissible range in (28), the algorithm diverges, causing $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|$ to increase gradually. This highlights the importance of maintaining an appropriate step size. Moreover, Figure 6 suggests that increasing $\gamma$ does not necessarily accelerate convergence. In fact, notice that the term $\gamma(2/\beta-\gamma)$ in (35) governs the nonexpansivity of $\mathcal{G}$; this reaches its maximum when $\gamma$ is neither too small nor too large.
@@ 5.4 Fusion Results
A comparison between HyDeFuse and various classical methods, including GSA [7], CNMF [17], GLP [13], MAPSMM [9], and HySure [8] is provided in Figure 9. The corresponding metrics are reported in Table 7. It is not surprising that HyDeFuse outperforms these classical methods.
We next compare HyDeFuse with two recent methods, GTTN [74] and CTDF [18], on the Pavia dataset. A visual comparison is provided in Figure 10, while various performance metrics are presented in Tables 5 and 6. We observe that our method, HyDeFuse, performs on par with these approaches. To further demonstrate its effectiveness and generalizability, we evaluate HyDeFuse under two different SNR settings—one used in the original papers of GTTN and CTDF, and another that we have consistently used for comparisons.
@@
Figure 9: Fusion results on the Pavia dataset. See Table 7 for the quality metrics.
@@
Figure 10: Comparison of recent tensor-based methods with Gaussian point spread function (model for $\mathbf{B}$) with radius $7$ and standard deviation 22. The corresponding metrics are shown in Table 5.
@@
CTDF 38.49 0.013 1.94 3.67 0.9869
GTNN $\mathbf{41.58}$ $\mathbf{0.008}$ $\mathbf{1.37}$ $\mathbf{2.39}$ $\mathbf{0.9859}$
@@
Table 6: Quality metrics with Gaussian point spread function (model for $\mathbf{B}$) with radius $7$ and standard deviation 22. In this experiment, $\text{SNR}_{h}=20\,\text{dB}$ and $\text{SNR}_{m}=20\,\text{dB}$.
Dataset Methods PSNR RMSE ERGAS SAM UIQI
Bicubic 25.05 0.0559 8.44 8.40 0.746
MASMM 28.63 0.0384 5.85 9.61 0.91
GLP 28.39 0.037 5.95 9.83 0.9018
Pavia CNMF 29.46 0.033 5.28 8.09 0.924
GSA 29.95 0.0318 5.03 9.98 0.918
HySure $\mathbf{34.12}$ $\mathbf{0.020}$ $\mathbf{3.029}$ $\mathbf{5.398}$ $\mathbf{0.965}$
Bicubic 23.05 0.0702 6.24 6.21 0.57
MASMM 25.01 0.0556 5.06 5.59 0.7761
GLP 24.36 0.059 5.43 6.35 0.751
Paris CNMF 25.39 0.05 4.85 4.55 0.80
GSA 24.48 0.06 5.36 5.78 0.78
HySure $\mathbf{26.82}$ $\mathbf{0.04}$ $\mathbf{4.13}$ $\mathbf{3.17}$ $\mathbf{0.839}$
Bicubic 29.02 0.035 7.49 7.30 0.54
MASMM 30.84 0.028 6.51 8.74 0.7170
GLP 30.20 0.030 6.80 8.31 0.698
Chikusei CNMF 34.14 0.019 3.97 5.23 0.847
GSA 32.29 0.020 6.89 8.84 0.761
Table 7: Comparison with classical methods for all three datasets.
(a) ground truth
(b) DnCNN
(c) BM4D
(d) GMM
(e) CasKD
Figure 11: PnP-PGD fusion on the Chikusei dataset where we compare our denoiser CasKD with other denoisers (as regularizers). The metrics are reported in Table 8.
Denoisers PSNR RMSE ERGAS SAM UIQI
DnCNN 42.29 0.0076 1.43 1.90 0.961
BM4D 42.40 0.0075 1.47 1.97 0.967
GMM 35.67 0.016 2.78 5.25 0.872
CasKD 42.27 0.008 1.39 1.79 0.967
Table 8: Quality metrics for Figure 11.
Finally, we analyze the impact of replacing our CasKD denoiser in HyDeFuse with existing denoisers, including DnCNN [39], BM4D [75], and GMM [45]. As shown in Figure 11, the reconstruction quality of CasKD is comparable to that of DnCNN and BM4D, and in some cases, it even surpasses them. The reported fused images were obtained after 100 iterations. We note that it is not expected that a simple kernel denoiser such as CasKD –– that essentially works by averaging local pixels –– can beat a trained denoiser such as DnCNN or a sophisticated denoiser such as BM4D. However, note that we are using CasKD as a regularizer not as a denoiser, where it is applied repeatedly across the iterations. This iterative regularization is what enables CasKD to achieve strong fusion results.
Dataset SNR Denoiser PSNR RMSE ERGAS SAM UIQI
$\mathcal{W}$ 33.33 0.021 3.04 5.19 0.960
Pavia 20dB CasKD 35.89 0.016 2.39 3.61 0.980
$\mathcal{V}$ 35.36 0.017 2.60 3.88 0.977
$\mathcal{W}$ 38.81 0.011 2.09 3.19 0.927
Chikusei 20dB CasKD 42.27 0.0076 1.39 1.79 0.967
$\mathcal{V}$ 42.17 0.0078 1.42 1.84 0.963
Table 9: Comparison of CasKD and its component denoisers as implicit regularizers in HyDeFuse.

### v2

@@ 4.4 Discussion
We now discuss how our analysis and results connect with earlier works that employ kernel denoisers to construct provably convergent PnP models [47, 42, 55, 62]. These studies interpret the kernel denoiser as the proximal operator of a convex regularizer, thereby linking the PnP framework to a convex optimization setting. The convergence of the corresponding objective function is then established using standard results from proximal algorithm theory [74].
In contrast, the CasKD denoiser in HyDeFuse is not necessarily self-adjoint, since $\mathcal{V}$ and $\mathcal{W}$ do not commute. This asymmetry implies that $\mathcal{D}$ cannot be expressed as the proximal operator of any convex regularizer, this follows from a classical result due to Moreau [76, 77]. Consequently, the approach in [47, 42, 55] does not directly apply here.
Moreover, we establish a stronger result, namely the convergence of the iterates themselves (not just the objective), and further provide a linear convergence rate. While our approach is inspired by [62], the specific loss function, vector space, operators, denoiser, and supporting results differ substantially in our setting. Therefore, the analysis in [62] cannot be directly extended to HyDeFuse.
@@ 5 Numerical Results
We validate our theoretical results and demonstrate HyDeFuse in practice. We present numerical results on global convergence and convergence rate. We compare the proposed CasKD denoiser with existing ones and benchmark HyDeFuse against both classical and state-of-the-art fusion methods.
@@ 5.1 Experimental Setup
We begin by describing the dataset, the parameters of the forward model, the quality metrics used for evaluation, and the denoising and fusion methods used for comparison.
@@
Model parameters. To test robustness under different conditions, we model the blurring operator for the HS image in two ways. The main setup uses the Starck–Murtagh filter, as in [8, 47]. Following [78], we consider a Gaussian point spread function (PSF) with a radius of 7 and a standard deviation of 2. The downsampling factors used to create the observed HS image are 4 for Pavia, 6 for Chikusei, and 3 for Paris. The spectral response matrix $\mathbf{R}$ is computed as described in [8]. Unless stated otherwise, the signal-to-noise ratio (SNR) is fixed at 20 dB for both the HS and MS images, denoted by $\mathrm{SNR}_{h}$ and $\mathrm{SNR}_{m}$. The HS images of Pavia, Chikusei, and Paris contain $L_{h}=93$, $128$, and $128$ bands, while the corresponding MS images have $L_{m}=4$, $8$, and $9$ bands.
Hyperparameters. Following [8], we assume that the ground-truth image lies in a low-dimensional subspace, with dimension $L_{s}=10$ for the Pavia and Paris datasets and $L_{s}=5$ for the Chikusei dataset. The patch size is set to $7\times 7$, the search window to $10\times 10$, and the number of clusters to $40$. The parameters $\sigma_{1}$ and $\sigma_{2}$ are chosen from the range $(5/255,25/255)$ and tuned for each dataset. In our experiments, we fix the number of iterations to 100 rather than using an explicit stopping rule. However, a stopping condition, such as $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|_{\mathbb{H}}/\|\mathbf{X}_{k}\|_{\mathbb{H}}<\varepsilon$, can be easily integrated into our framework. The theoretical convergence of the iterative scheme guarantees that this condition will always be met for any chosen tolerance $\varepsilon$.
Compared methods. We compare our CasKD denoiser with both traditional and recent denoisiners. Specifically, we benchmark CasKD against BM4D [79], DnCNN [39], FHDD-NLM [47], KBR [46], LLRT [45], FastHyMix [49], and NGMeet [48]. For image fusion, we evaluate HyDeFuse against a range of classical and state-of-the-art methods, including GSA [7], CNMF [17], GLP [13], MAPSMM [9], HySure [8], GTTN [78], and CTDF [18].
Quality metrics. We employ standard quality metrics [8, 43, 78], including Peak Signal-to-Noise Ratio (PSNR), Root Mean Squared Error (RMSE), Spectral Angle Mapper (SAM), Erreur Relative Globale Adimensionnelle de Synthèse (ERGAS), and Universal Image Quality Index (UIQI). For reconstruction tasks, lower values of RMSE, SAM, and ERGAS indicate better performance, whereas higher PSNR and UIQI values (with a maximum of $1$) correspond to superior reconstruction quality.
@@ 5.2 Denoising Results
The CasKD denoiser combines two kernel denoisers. While combining denoisers need not always improve performance, the two components play different roles in CasKD—they handle inter- and intra-band correlations between patches. This likely explains the improved results seen in Table 1.
To further assess CasKD, we compare it with several well-known denoisers: BM4D [79], DnCNN [39], FHDD-NLM [47], KBR [46], LLRT [45], FastHyMix [49], and NGMeet [48]. A visual comparison is presented in Figure 4. A small region is marked with a red box, and its zoomed-in view is shown in the lower-right corner to help access the denoising quality. Because the noise is relatively low ($\sigma=5/255$), most denoisers produce visually good results; however, the insets make it easier to observe subtle texture differences among the methods.
Table 2 provides a detailed quantitative comparison using PSNR and UIQI. CasKD performs competitively with existing denoisers and, in some cases, achieves the best results.
Methods PSNR (5/255) UIQI (5/255) PSNR (20/255) UIQI (5/255)
@@
Table 2: Comparison of denoising results (with Gaussian noise) on the Pavia dataset for zero-mean Gaussian noise with standard deviations of 5/255 and 20/255 (shown in brackets). The top two values are in bold, and the best value is underlined. This convention is used in all the tables.
@@
Figure 4: Visual comparison of CasKD with other denoisers on the Pavia dataset. Zero-mean Gaussian noise with a standard deviation of $5/255$ is independently added to pixels across all bands. For visualization, three selected bands are shown.
@@ 5.3 Convergence Results
We evaluate the convergence behavior of HyDeFuse under various settings. First, we analyze the empirical convergence rate and present evidence of its contractive nature. Next, we examine the effect of the step size on convergence and its relation to linear convergence. We then assess the stability of the reconstruction in terms of PSNR for the same experiments. Finally, we test the uniqueness of the fixed point of HyDeFuse.
We first estimate the contraction factor $\mu$ in (24), which, as shown in (30), determines the convergence rate of HyDeFuse. We use the power method [70] to compute $\mu$, the largest singular value of the linear operator $\mathcal{P}$. A direct computation using SVD is impractical, since $\mathcal{P}$ is too large to store explicitly. Instead, the power method approximates $\mu$ by iteratively applying $\mathcal{P}$ and its adjoint, which is given by $\mathcal{P}^{*}=\mathcal{G}\circ\mathcal{W}\circ\mathcal{V}$, since $\mathcal{V}$, $\mathcal{W}$, and $\mathcal{G}$ are self-adjoint.
Figure 5: Variation of the convergence rate $\mu$ for different values of $\lambda$ in the loss function (6), evaluated on the Chikusei dataset. The parameters $\sigma_{1}$ and $\sigma_{2}$ (reported without scaling by $1/255$) represent the standard deviations of the Gaussian (RBF) kernels used in $\mathcal{W}$ and $\mathcal{V}$, respectively.
@@
Table 3: Contraction factors are shown for different step sizes and denoiser settings. These results are evaluated on the Chikusei dataset. The step size $\gamma$ is expressed in multiples of $1/\beta$, where $\beta$, computed using the power method, is found to be $0.9757$. The pair $(\sigma_{1},\sigma_{2})$ represents the standard deviations of the Gaussian (RBF) kernels used in $\mathcal{W}$ and $\mathcal{V}$, with $a=5/255$, $b=10/255$, and $c=30/255$.
In Table 3, we validate the contraction factor $\mu$ for HyDeFuse using the Chikusei dataset. The value of $\mu$ depends on the step size $\gamma$, the forward model parameters $\mathbf{A}$ and $\mathbf{R}$, and the denoiser $\mathcal{D}$. We compute $\mu$ using the power method for step sizes $0<\gamma<2/\beta$ and for different denoiser settings. As predicted by Theorem 1, the contraction factor remains below $1$.
We also compute $\mu$ for different values of $\lambda$, which controls the balance in the loss function. The results, shown in Figure 5, again confirm that $\mu<1$. As seen in the plot, $\mu$ approaches $1$ as $\lambda$ tends to 0, but remains slightly below it, with $\mu\approx 1-\varepsilon$ where $\varepsilon\approx 10^{-3}$. Since these values are very close to $1$, a mathematical proof is important to rule out numerical artifacts that could make $\mu$ appear smaller than $1$.
Additionally, we see that convergence is faster when $\lambda$ is neither too large nor too small, which is noteworthy as this range is also expected to yield optimal fusion performance. For the Chikusei dataset (Figure 5), optimal fusion performance is observed for $\lambda$ in the range 4–8. This is consistent with the observation that intermediate $\lambda$ values yield optimal convergence. A similar trend is seen in the Pavia and Paris datasets.
Figure 6: The plot shows the successive differences between HyDeFuse iterates and their corresponding PSNR values for different step sizes $\gamma$ (in units of $1/\beta$) on the Paris dataset. Using the power method, we find $\beta=0.5998$ for this setup. Convergence is achieved for $\gamma\in(0,2/\beta)$, but as seen with $\gamma=2.2/\beta$, it fails outside this range. The PSNR stabilizes within about 10 iterations for an optimal choice of $\gamma$.
(a) ground truth
(b) $\mathbf{X}_{0}$ : all-ones
(c) $\mathbf{X}_{0}$ : all-zeros
(d) $\mathbf{X}_{0}$ : white noise
Figure 7: HyDeFuse results on the Chikusei dataset for different initializations $\mathbf{X}_{0}$. In all cases, the performance metrics remain identical (PSNR = 42.27, RMSE = 0.0076, ERGAS = 1.39, SAM = 1.79, and UIQI = 0.967), confirming that the HyDeFuse iterates converge to the same fixed point.
We next study the convergence of the iterates $\{\mathbf{X}_{k}\}$ generated using (9). Figure 6 shows a sample result on the Paris dataset using the Starck–Murtagh filter as the blurring operator. As predicted by Theorem 1, the successive difference $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|_{\mathbb{H}}$ decreases to zero, and the PSNR stabilizes within about 10–15 iterations a trend consistently observed across all datasets.
When the step size $\gamma$ exceeds the admissible range $0<\gamma<2/\beta$, we see in Figure 6 that the algorithm diverges, with $\|\mathbf{X}_{k}-\mathbf{X}_{k-1}\|_{\mathbb{H}}$ gradually increasing. This shows the importance of choosing $\gamma$ properly. Moreover, we notice that increasing $\gamma$ does not necessarily speed up convergence.
@@
To further demonstrate the contractive nature of the reconstruction operator $\mathcal{T}$, we perform fusion experiments on the Chikusei dataset with different initializations $\mathbf{X}_{0}$. As shown in Figure 7, all runs converge to the same reconstruction, confirming the uniqueness of the fixed point of the contractive operator $\mathcal{T}$.
The results in this subsection provide clear experimental evidence of the convergence of HyDeFuse. Using the power method, we show that the reconstruction operator $\mathcal{P}$ is contractive across different settings, with the contraction factor $\mu$ always less than $1$. This holds for various step sizes $\gamma$ and values of the balancing parameter $\lambda$, consistent with Theorem 1. The iterates exhibit linear convergence, and the PSNR stabilizes within 10–15 iterations. We also find that large values of $\gamma$ can lead to divergence, emphasizing the need for proper parameter selection. Finally, the fixed-point analysis confirms that HyDeFuse consistently converges to a unique solution from different initializations.
@@ 5.4 Fusion Results
Finally, we compare the performance of HyDeFuse with several classical methods as well as recent state-of-the-art techniques, under two different noise levels.
A comparison between HyDeFuse and various classical methods, including GSA [7], CNMF [17], GLP [13], MAPSMM [9], and HySure [8] is provided in Figure 8. In this figure, a small region of the original image is marked with a red box, and a zoomed-in view is shown in the lower-right corner. This close-up comparison clearly shows that HyDeFuse performs better than the other methods. The corresponding metrics are reported in Table 4.
We next compare HyDeFuse with two recent methods: GTTN [78] and CTDF [18]. The experiments are conducted on the Pavia dataset. A visual comparison is provided in Figure 9, while various performance metrics are presented in Tables 5 and 6. We observe that our method, HyDeFuse, performs on par with these approaches.
To further test its effectiveness and generality, we evaluate HyDeFuse under two different SNR settings—one used in the original GTTN and CTDF papers, and another used consistently in our comparisons. Under high noise in both the HS and MS images, HyDeFuse clearly outperforms these recent methods, as shown in Table 6. This improvement is mainly due to the robustness of the CasKD denoiser, which adapts well to different noise levels. Some differences in performance may also be due to hyperparameter tuning, which could potentially improve the performance of competing methods.
In summary, HyDeFuse consistently outperforms classical fusion methods. Compared to recent methods such as GTTN and CTDF, it remains competitive under standard conditions and achieves significant improvements at higher noise levels, due to the adaptive capabilities of the CasKD denoiser.
@@
Figure 8: Fusion results on the Pavia dataset. The quality metrics are shown in Table 4.
@@
Figure 9: Comparison of recent tensor-based methods on the Pavia dataset. The blur operator $\mathbf{B}$ is modeled using a Gaussian point spread function (PSF) with a radius of 7 and a standard deviation of 2. The corresponding performance metrics are listed in Table 5.
Dataset Methods PSNR RMSE ERGAS SAM UIQI
Bicubic 25.05 0.0559 8.44 8.40 0.746
MASMM 28.63 0.0384 5.85 9.61 0.910
GLP 28.39 0.037 5.95 9.83 0.901
Pavia CNMF 29.46 0.033 5.28 8.09 0.924
GSA 29.95 0.0318 5.03 9.98 0.918
HySure $\mathbf{34.12}$ $\mathbf{0.020}$ $\mathbf{3.02}$ $\mathbf{5.39}$ $\mathbf{0.965}$
Bicubic 23.05 0.0702 6.24 6.21 0.57
MASMM 25.01 0.0556 5.06 5.59 0.7761
GLP 24.36 0.059 5.43 6.35 0.751
Paris CNMF 25.39 0.05 4.85 4.55 0.80
GSA 24.48 0.06 5.36 5.78 0.78
HySure $\mathbf{26.82}$ $\underline{\mathbf{0.04}}$ $\mathbf{4.13}$ $\mathbf{3.17}$ $\mathbf{0.839}$
Bicubic 29.02 0.035 7.49 7.30 0.54
MASMM 30.84 0.028 6.51 8.74 0.717
GLP 30.20 0.030 6.80 8.31 0.698
Chikusei CNMF 34.14 0.019 3.97 5.23 0.847
GSA 32.29 0.020 6.89 8.84 0.761
Table 4: Comparison of HyDeFuse with classical fusion methods across three datasets.
@@
CTDF 38.49 0.013 1.94 3.67 $\mathbf{0.9869}$
GTNN $\mathbf{41.58}$ $\mathbf{0.008}$ $\mathbf{1.37}$ $\mathbf{2.39}$ $0.9859$
@@
Table 6: Quality metrics evaluated on the Pavia dataset. The blurring operator $\mathbf{B}$ is modeled using a Gaussian point spread function (PSF) with a radius of 7 and a standard deviation of 2. In this experiment, $\text{SNR}_{h}=20\,\text{dB}$ and $\text{SNR}_{m}=20\,\text{dB}$.